# Chaotic stochastic resonance in Mackey–Glass equations

### v3

@@ III.3 Resonant period and unstable spiral
The classical approximation, $T\simeq\tau$, holds only when $\tau$ is extremely large. Numerical validation confirms that the primary resonant frequency, $f_{1}^{*}$ (black points), agrees well with the theoretical estimate, $s_{2}$ (red line), as shown in Figs. 5 (b) and (d).
Figure 5: Unstable spirals in the weakly and strongly nonlinear regimes. Phase-space portraits projected onto $(x(t),x(t-\tau))$ for (a) $a=a_{1}$ and (c) $a=a_{2}$. Transients initialized near the origin (red dots) eventually converge to other stable attractors (blue dots). The trajectory is computed from the initial function $\phi(t)=0.01\sin(2\pi f^{*}_{1}t)+0.02\sin(2\pi f^{*}_{2}t)$ and departs from the origin following a spiral structure. The primary resonant frequency $f_{1}^{*}$ as a function of $\tau$ for (b) $a=a_{1}$ and (d) $a=a_{2}$. The theoretical estimate (4) (red lines) agrees well with numerical results (black points). The classical estimate $f_{1}^{*}=1/\tau$ is also shown (dotted lines).
@@ IV Conclusion
We investigated the dynamics of the Mackey–Glass equation in the presence of noise. In the weakly nonlinear regime, stable SR emerges as switching dynamics between two quasi-stationary states, with the largest Lyapunov exponent remaining negative at the resonance point. In this regime, the resonance point $\sigma^{*}$ follows the zero-crossing point of the largest Lyapunov exponent $\sigma_{0}$ ($\sigma_{0}<\sigma^{*}$). The deterministic chaotic attractor becomes stabilized into a random point attractor, and a resonance phenomenon in the classical double well potential model Benzi, Sutera, and Vulpiani (1981) effectively occurs. In the strongly nonlinear regime, we discovered chaotic SR, characterized by the coexistence of SR and stochastic chaos with positive Lyapunov exponents. Here, the resonance point $\sigma^{*}$ precedes the zero-crossing point $\sigma_{0}$ ($\sigma^{*}<\sigma_{0}$), indicating that the resonance occurs while the system remains chaotic. We also numerically observed the existence of chaotic SR in the stochastic Duffing equation. Under the overdamped limit, the stochastic Duffing equation reduces to the classical double well potential model, which exhibits stable SR. Thus, chaotic SR is expected to represent a universal class of noise-induced phenomena in strongly nonlinear random dynamical systems. Further investigation of noise-induced behavior in generic multi-attractor systems represents a promising avenue for future research.
@@ Acknowledgements.
Authors thank Prof. T. Ohira (Nagoya University) for valuable comments. Y.S. is supported by JSPS Grant-in-Aid for Scientific Research (B), JP No. 21H01002.
@@ Appendix A Definitions of random pullback attractor and random Lyapunov exponent
To parametrize a noise realization, $\omega\in\Omega$, in time $t$, we introduce a family of measure-preserving maps, $\theta_{t}:\varOmega\rightarrow\varOmega$ satisfied with $\theta_{0}=\rm{id}_{\Omega}$ and $\theta_{s+t}=\theta_{s}\circ\theta_{t}$ for all $t$. The time evolution of random dynamical systems is described by a stochastic flow, $\varPhi(t,\omega):X\rightarrow X$, satisfying the cocycle property, $\varPhi(t+s,\omega)=\varPhi(t,\theta_{s}(\omega))\circ\varPhi(s,\omega)$. For systems with a single attractor, a set $\mathcal{A}(\omega)$ is defined as a random pullback attractor if it satisfies the following three conditions (see Arnold (1998); Chekroun, Simonnet, and Ghil (2011) for the exact definition):
@@
    $\mathcal{A}(\omega)$ is a compact, i.e., $A(\omega):=\{x\in X|~(\omega,x)\in\mathcal{A}(\omega)\}\subset X$ is compact for almost all $\omega\in\Omega$.
@@
    $\mathcal{A(\omega)}$ is pullback attracting, i.e., $\lim_{t\rightarrow\infty}d_{X}(\varPhi(t,\theta_{-t}\omega)B,A(\omega))=0$ holds for all $B\subset X$ and for almost all $\omega\in\Omega$.
@@
A realization $A(\omega)$ of a random pullback attractor ${\mathcal{A}}(\omega)$ can be numerically approximated by a snapshot of trajectories evolving from a set of initial values $B$, with a fixed noise realization $\omega$, and with a integration period given by the pullback time $t_{p}$. For a precise numerical computation, a large number of initial points for $B$ and a large pullback time $t_{p}$ are adopted. The Lyapunov exponent of stochastic dynamics on random pullback attractors is given by the average expansion rate of perturbations, similarly to those of deterministic dynamics. For example, the Lyapunov exponent of a one-dimensional stochastic differential equation $dx=f(x)dt+\sigma dW_{t}$, where $W_{t}$ is the Wiener process, is given by
@@
where $\varPhi$ represents the stochastic flow, and $x_{0}\in X$ is the initial condition. Note that the Lyapunov exponent is a random variable in general. When the system is ergodic and has a single attractor, $\lambda(\omega,x_{0})$ is a constant for almost all $\omega$ and for all $x_{0}$. As for high-dimensional random dynamical systems, the Lyapunov spectrum is defined in a similar manner to those of high-dimensional deterministic dynamical systems.
@@ Appendix B Bifurcation diagram of random pullback attractors
We construct a bifurcation diagram of pullback attractors for the SMG. Plotting trajectories starting from many initial conditions with a fixed noise realization for each parameter, the bifurcation of random pullback attractors is successfully visualized. Similarly to random pullback attractors, this bifurcation diagram also evolves in time. In Fig. 6, emergence of the pseudo-periodic motion of the measure of all possible trajectories $\rho_{t}$ is observed near $a=0.14$ and $\sigma=0.15$.
Figure 6: Snapshots of bifurcation diagrams of random pullback attractors with a fixed noise realization for $\sigma=0.15$ and varying $a$. For all parameters, random pullback attractors are computed using a fixed noise realization. The numerical scheme follows that used in Fig. 1.
@@ Appendix C Linear mode analysis of unstable spirals
When the delay $\tau$ is sufficiently large, the peaks of the power spectra of the Mackey–Glass equation (MG) agree with the linear modes around fixed points Mensour and Longtin (1998). The characteristic equations at fixed points $x=0,\pm x^{*}$, where $x^{*}=(\frac{a-b}{b})^{1/c}$
@@
where $W_{n}(z)$ is the Lambert W-function Corless et al. (1996). Using the asymptotic expansion of the Lambert W-function Amann, Schöll, and Just (2007), the frequencies of linear modes around the fixed points are given by the imaginary part of $\lambda_{n},\mu_{n}$, respectively expressed as
@@
The even modes $s_{2n}$ represent an unstable spiral around the origin, and the odd modes $s_{2n+1}$ represent spirals around the fixed point $x=\pm x^{*}$. The characteristic frequencies of the dynamics dominated by odd modes at $x=\pm x^{*}$ (Fig. 7). Adding the optimal noise, the odd modes are suppressed and the even modes become dominant, causing the emergence of stochastic resonance (SR) with resonant frequencies corresponding to the even mode.
Figure 7: (a) Roots of linear modes $\chi=\lambda_{n}$ (black) and $chi=\mu_{n}$ (blue) at $a=a_{1}=0.14$. (b) Power spectra for $a=a_{1}=0.14$ with $\sigma=0$ (blue line) and optimal noise $\sigma=\sigma^{*}$ (black line).

### v4

@@ III.3 Resonant period and unstable spiral
The classical approximation, $T\simeq\tau$, holds only when $\tau$ is extremely large. Numerical validation confirms that the primary resonant frequency, $f_{1}^{*}$ (black points), agrees well with the theoretical estimate, $s_{2}$ (red line), as shown in Figs. 5 (b) and (d). This result suggests that SR is associated with frequent visits to the unstable spiral caused by noise. Since our approximation is based on these spiral structures, it is applicable to a broad class of DDEs beyond the specific case of the Mackey–Glass equation.
Figure 5: Unstable spirals in the weakly and strongly nonlinear regimes. Phase-space portraits projected onto $(x(t),x(t-\tau))$ plane for (a) $a=a_{1}$ and (c) $a=a_{2}$. Transients initialized near the origin (red dots) eventually converge to other stable attractors (blue dots). The trajectory is computed from the initial function $\phi(t)=0.01\sin(2\pi f^{*}_{1}t)+0.02\sin(2\pi f^{*}_{2}t)$ and departs from the origin following a spiral structure. The primary resonant frequency $f_{1}^{*}$ as a function of $\tau$ for (b) $a=a_{1}$ and (d) $a=a_{2}$. The theoretical estimate, the reciprocal of Eq. (4), (red lines) agrees well with numerical results (black points). The classical estimate $f_{1}^{*}=1/\tau$ is also shown (dotted lines).
@@ IV Conclusion
We investigated the dynamics of the Mackey–Glass equation in the presence of noise. In the weakly nonlinear regime, stable SR emerges as switching dynamics between two quasi-stationary states, with the largest LE remaining negative at the resonance point. In this regime, the resonance point $\sigma^{*}$ follows the zero-crossing point of the largest LE $\sigma_{0}$ ($\sigma_{0}<\sigma^{*}$). The deterministic chaotic attractor becomes stabilized into a random point attractor, and a resonance phenomenon analogous to that in the classical double-well potential model Benzi, Sutera, and Vulpiani (1981) effectively occurs. In the strongly nonlinear regime, we identify chaotic SR, characterized by the coexistence of SR and stochastic chaos with positive LEs. Here, the resonance point $\sigma^{*}$ precedes the zero-crossing point $\sigma_{0}$ ($\sigma^{*}<\sigma_{0}$), indicating that resonance occurs while the system remains chaotic.
We observe chaotic SR in the Mackey-Glass equation as well as chaotic SR in the Duffing equation and the underdamped FitzHugh-Nagumo equation (see Appendix-D), demonstrating the universality of chaotic SR across a broad class of strongly nonlinear random dynamical systems.
@@ Acknowledgements.
Authors thank Prof. T. Ohira (Nagoya University), Prof. Hiroya Nakao (Institute of Science Tokyo), and Prof. Hiroshi Kori (University of Tokyo) for valuable comments. Authors are supported by JSPS Moonshot project No. JPMJMS2282-15. Y.S. is supported by JSPS Grant-in-Aid for Scientific Research (B), JP No. 21H01002.
@@ Appendix A Definitions of random pullback attractor and random Lyapunov exponent
To parametrize a noise realization $\omega\in\Omega$ in time $t$, we introduce a family of measure-preserving maps $\theta_{t}:\varOmega\rightarrow\varOmega$ satisfied with $\theta_{0}=\rm{id}_{\Omega}$ and $\theta_{s+t}=\theta_{s}\circ\theta_{t}$ for all $t$. The time evolution of random dynamical systems is described by a stochastic flow $\varPhi(t,\omega):X\rightarrow X$ satisfying the cocycle property $\varPhi(t+s,\omega)=\varPhi(t,\theta_{s}(\omega))\circ\varPhi(s,\omega)$. For systems with a single attractor, a set $\mathcal{A}(\omega)$ is defined as a random pullback attractor if it satisfies the following three conditions (see Arnold (1998); Chekroun, Simonnet, and Ghil (2011) for the exact definition):
@@
    $\mathcal{A}(\omega)$ is compact, i.e., $A(\omega):=\{x\in X|~(\omega,x)\in\mathcal{A}(\omega)\}\subset X$ is compact for almost all $\omega\in\Omega$.
@@
    $\mathcal{A(\omega)}$ is attracting in the pullback sense, i.e., $\lim_{t\rightarrow\infty}d_{X}(\varPhi(t,\theta_{-t}\omega)B,A(\omega))=0$ holds for all $B\subset X$ and for almost all $\omega\in\Omega$.
@@
A realization $A(\omega)$ of a random pullback attractor ${\mathcal{A}}(\omega)$ can be numerically approximated by a snapshot of trajectories evolving from a set of initial values $B$, under a fixed noise realization $\omega$, for a integration period given by the pullback time $t_{p}$. For a precise numerical computation, a large number of initial points for $B$ and a sufficiently long pullback time $t_{p}$ are adopted. Lyapunov exponents (LEs) for stochastic dynamics on random pullback attractors are defined as the average expansion rates of infinitesimal perturbations, in direct analogy with deterministic systems. As a simple example, the LE of a one-dimensional stochastic differential equation $dx=f(x)dt+\sigma dW_{t}$, where $W_{t}$ denotes a Wiener process, is given by
@@
where $\varPhi$ denotes the stochastic flow and $x_{0}\in X$ is the initial state. In general, $\lambda(\omega,x_{0})$ is a random variable. For ergodic systems with a single attractor, $\lambda(\omega,x_{0})$ takes constant value for almost all $\omega$ and for any $x_{0}$. For high-dimensional random dynamical systems, the Lyapunov spectrum is defined in a manner analogous to that of high-dimensional deterministic dynamical systems.
@@ Appendix B Bifurcation diagram of random pullback attractors
We construct a bifurcation diagram of random pullback attractors for the SMG. Plotting trajectories starting from many initial conditions under a fixed noise realization for each value of the parameter $a$, the bifurcation of random pullback attractors is successfully visualized. Similar to the attractors themselves, the bifurcation diagram evolves in time. In Fig. 6, the pseudo-periodic oscillation of the distribution $\rho_{t}$ of all possible trajectories appears near $a=0.14$ and $\sigma=0.15$, illustrating the onset of SR behavior.
Figure 6: Snapshots of bifurcation diagrams of random pullback attractors for $\sigma=0.15$ as the parameter $a$ is varied under a fixed noise realization. For all parameters, random pullback attractors are computed using a fixed pullback time $t_{p}=4\times 10^{4}$. The numerical scheme follows that used in Fig. 1.
@@ Appendix C Linear mode analysis of unstable spirals
When the delay $\tau$ is sufficiently large, the peaks in the power spectra of the MG correspond to the linear modes around fixed points Mensour and Longtin (1998). The characteristic equations at fixed points $x=0,\pm x^{*}$, where $x^{*}=(\frac{a-b}{b})^{1/c}$,
@@
where $W_{n}(z)$ denotes the Lambert W-function Corless et al. (1996). Using the asymptotic expansion of the Lambert W-function Amann, Schöll, and Just (2007), the frequencies associated with the linear modes are given by the imaginary parts of $\lambda_{n}$ and $\mu_{n}$:
@@
The even modes $s_{2n}$ describe the unstable spiral around the origin, while the odd modes $s_{2n+1}$ characterize spirals around $x=\pm x^{*}$. The deterministic dynamics dominated by odd modes. When the noise is added, the odd modes are suppressed and the even modes become dominant (Fig. 7), producing SR with resonant frequencies associated with the even linear modes.
Figure 7: (a) Roots of linear modes $\chi=\lambda_{n}$ (black) and $chi=\mu_{n}$ (blue) at $a=a_{1}=0.14$. (b) Power spectra for $a=a_{1}=0.14$ with $\sigma=0$ (blue) and optimal noise $\sigma=\sigma^{*}$ (black).
Appendix D Universality of chaotic stochastic resonance
Chaotic SR arises not only in time-delayed systems exhibiting coherence resonance, but also in periodically forced systems exhibiting classical SR and in excitable systems exhibiting coherence resonance.
To illustrate classical SR, we consider an overdamped dynamics in a double-well potential,
$dx=(x-x^{3})ds.$ (13)
Upon adding a slow periodic external force and Gaussian noise, we obtain
$dx=\left[x-x^{3}+\varepsilon\cos(\Omega s)\right]ds+\sigma^{\prime}dW_{s},$ (14)
where $\sigma^{\prime}$ represents the noise intensity, and $W_{s}$ denotes the Wiener process. The external force is specified by the amplitude $\varepsilon$ and the angular frequency $\Omega$. For $\varepsilon=0.1$ and $\Omega=2\pi/100$, we observe stable SR in Eq. (14) at $\sigma^{\prime}\simeq 0.354$ Gammaitoni et al. (1998). Next, we consider the underdamped dynamics
$dx$ $=$ $ydt$
$dy$ $=$ $(-\gamma y+x-x^{3})dt,$ (15)
where $\gamma$ is the damping rate. In the overdamped limit ($\gamma\rightarrow\infty$, $t=\gamma s$, and $dy/dt\rightarrow 0$), Eq. (15) reduces to Eq. (13). By adding a slow periodic external force and Gaussian noise, we obtain the stochastic Duffing equation given by
$dx$ $=$ $ydt$
$dy$ $=$ $\left[-\gamma y+x-x^{3}+\varepsilon\cos(\omega t)\right]dt+\sigma dW_{t},$ (16)
where $\sigma$ represents the noise intensity, and $W_{t}$ denotes the Wiener process. The external force is specified by the amplitude $\varepsilon$ and the angular frequency $\omega$. For $\gamma=0.2$, $\varepsilon=0.3$, and $\omega=0.1$, we observe chaotic SR in Eq. (16) at $\sigma\simeq 0.22$ (see FIG.8). Similar phenomena have been reported in both numerical and laboratory experiments Wiebe and Spottswood (2014).
Figure 8: Chaotic SR in the Duffing equation with $\gamma=0.2$, $\varepsilon=0.3$, and $\omega=0.1$. (a) Power at the resonant frequency $S(f^{*})$, $f^{*}\simeq 0.0159$ (black points), and the largest random LE $\lambda_{1}$ (red points) as functions of $\sigma$. At the resonance point $\sigma=\sigma^{*}\simeq 0.22$, the largest random LE $\lambda_{1}\simeq 0.018>0$. (b) Snapshot of the random pullback attractor at the resonance point, computed from $2\times 10^{4}$ initial conditions uniformly distributed over $2,2]\times[-2,2]$ with a pullback time $t_{p}=400$.
For coherence resonance, we consider the FitzHugh-Nagumo equation,
$\varepsilon dx$ $=$ $(x-\frac{x^{3}}{3}-y)ds,$
$dy$ $=$ $(x+a)ds$ (17)
where $\varepsilon$ represents the timescale separation between $x$ and $y$ and $a$ is the excitation threshold. With additive Gaussian noise, we obtain
$\varepsilon dx$ $=$ $(x-\frac{x^{3}}{3}-y)ds,$
$dy$ $=$ $(x+a)ds+\sigma^{\prime}dW_{s}$ (18)
where $\sigma^{\prime}$ represents the noise intensity, and $W_{s}$ denotes the Wiener process. For $\varepsilon=0.01$ and $a=1.05$, we observe stable SR in Eq. (18) at $\sigma^{\prime}\simeq 0.07$ Pikovsky and Kurths (1997). Next we consider the underdamped FitzHugh-Nagumo equation Scott (1975)
$dx$ $=$ $zdt,$
$dy$ $=$ $\frac{\varepsilon}{\gamma}(x+a)dt$
$dz$ $=$ $(x-\frac{x^{3}}{3}-y-\gamma z)dt,$ (19)
where $\gamma$ is the damping rate, and the overdamped limit ($\gamma\rightarrow\infty$, $t=\gamma s/\varepsilon$, and $dz/dt\rightarrow 0$) reduces to (17). With additive Gaussian noise, we obtain
$dx$ $=$ $zdt,$
$dy$ $=$ $\frac{\varepsilon}{\gamma}(x+a)dt+\sigma dW_{t}$
$dz$ $=$ $(x-\frac{x^{3}}{3}-y-\gamma z)dt,$ (20)
where $\sigma$ represents the noise intensity, and $W_{t}$ denotes the Wiener process. For $\gamma=0.41$, $\varepsilon=0.1$, and $a=1.1$, we observe chaotic SR in Eq. (20) at $\sigma\simeq 0.12$ (see Fig. 9).
Figure 9: Chaotic SR in the underdamped FitzHugh–Nagumo equation with $\gamma=0.41$, $\varepsilon=0.1$, and $a=1.1$. (a) Power at the resonant frequency $S(f^{*})$, $f^{*}\simeq 0.037$ (black points), and the largest LE $\lambda_{1}$ (red points) as functions of $\sigma$. At the resonance point $\sigma=\sigma^{*}\simeq 0.12$, the Lyapunov exponent is $\lambda\simeq 0.003>0$. (b) Snapshot of the random pullback attractor, projected onto $(x(t),y(t))$ plane, at the resonance point $\sigma=\sigma^{*}$, computed from $2\times 10^{4}$ initial conditions uniformly distributed over $2,2]\times[-2,2]\times[-2,2]$ with a pullback time $t_{p}=400$.